# ORCID 0000-0001-6130-9413

### 03 Mar 2026 v1

@@ 1. Introduction
where $(i)$ denotes the period in the ss–adic representation of the number. We call such numbers ss–adic rational. To define the $n$th digit $\alpha_{n}(x)$ of a number $x$ as a function of $x$, we agree to use only the first ss–adic representation, namely, the one that has period $(0)$.
@@
The concept of the asymptotic mean of digits of a number is closely related to the concept of digit frequency. In the case of the binary numeral system, these concepts coincide. Let us recall this notion.
Let $N_{i}(x,n)$ denote the quantity of digits $i\in\mathcal{A}s$ in the ss–adic representation $\Delta^{s}{\alpha_{1}\alpha_{2}\ldots\alpha_{k}\ldots}$ of a number $x\in[0,1]$ up to and including the $n$th place, that is,
@@
We call a number $x$ normal in base $s$ (weakly normal) if for each $i\in\mathcal{A}_{s}$ the frequency exists and equals $\nu_{i}(x)=s^{-1}$.
@@
We focus on the case $s=4$, since the case $s=3$ is analyzed in detail in our paper [10]. In fact, for $s>3$, the set $S_{\theta}$ exhibits richer properties.
@@ 2. The set $\Theta_{1}$ and the Besicovitch–Eggleston sets
If $\theta=0$ or $\theta=3$, then $\Theta_{1}$ is an anomalously fractal set that is dense everywhere.
@@
Let $\theta=0$. If $\lim\limits_{n\to\infty}r_{n}(x)=0$ then for each $i\in{1,2,3}$ the following condition holds: $0\leqslant v^{(i)}_{n}(x)\leqslant v^{(1)}_{n}(x)+2v^{(2)}_{n}(x)+3v^{(3)}_{n}(x)=r_{n}(x)\to 0$ when $n\to\infty$, hence $\nu_{i}(x)=\lim\limits_{n\to\infty}v^{(i)}_{n}(x)=0$ and respectively $\nu_{0}(x)=1$. Therefore, $S_{\theta}=\Theta_{1}=E[1,0,0,0]$. This set is everywhere dense set and its Hausdorff–Besicovitch dimension is equal to
@@
Now let $\theta=3$. If $\lim\limits_{n\to\infty}r_{n}(x)=3$, then multiplying the first equation of system ()(*) by 3 and subtracting the second equation of the system, we obtain: $3v^{(0)}_{n}+2v^{(1)}_{n}+v^{(2)}_{n}=3-r_{n}$. Hence, $0\leqslant v^{(i)}_{n}(x)\leqslant 3v^{(0)}_{n}(x)+2v^{(1)}_{n}(x)+v^{(2)}_{n}(x)=3-r_{n}(x)\to 0$ when $n\to\infty$. Hence, $\nu_{i}(x)=0$ for all $i\in\{0,1,2\}$ and $\nu_{3}(x)=1$. Therefore, $\Theta_{2}=\Theta_{3}=\varnothing$ and $S_{\theta}=\Theta_{1}=E[0,0,0,1]$. This set is everywhere dense, and its Hausdorff–Besicovitch dimension equals 0. ∎
It is clear that if the 44–adic representation of a number $x$ has frequencies of all digits $\nu_{0}$, $\nu_{1}$, $\nu_{2}$, $\nu_{3}$, then it has an asymptotic mean of digits $r(x)$, given by
@@
If $\theta\neq\displaystyle\frac{3}{2}$, then $\Theta_{1}$ contains no normal numbers. Since almost all numbers (in the sense of Lebesgue measure) are normal, we have $\lambda(\Theta_{1})=0$. On the other hand, for $\theta=\displaystyle\frac{3}{2}$, the following inclusion holds: $\Theta_{1}\supset E\left[\displaystyle\frac{1}{4},\displaystyle\frac{1}{4},\displaystyle\frac{1}{4},\displaystyle\frac{1}{4}\right]$, where $\lambda(E)=1$. Thus, $\lambda(\Theta_{1})=1$.
@@
Оскiльки $E[p_{0},p_{1},p_{2},p_{3}]\subset\Theta_{1}$, то $\alpha_{0}(\Theta_{1})\geqslant\alpha_{0}(E[p_{0},p_{1},p_{2},p_{3}])=-\displaystyle\frac{m(\theta)}{\ln 4}.$
@@ 3. An example of a number from the set $\Theta_{1}$
We construct the number $x$ as follows. In the first step, we sequentially write $\tau^{\prime}{01}$ zeros, $\tau^{\prime}{11}$ ones, $\tau^{\prime}{21}$ twos, and $\tau^{\prime}{31}$ threes. After the kk‑th step, in the (k+1)(k+1)‑th step we append to the already written sequence $\tau^{\prime}{0k}$ zeros, $\tau^{\prime}{1k}$ ones, $\tau^{\prime}{2k}$ twos, and $\tau^{\prime}{3k}$ threes. As a result, among the first $\sum_{i=0}^{3}\tau_{in}$ symbols of the quaternary representation of $x$, there are exactly $\tau_{in}$ digits equal to ii. Let $n$ be a sufficiently large natural number. Since for any $x\in\mathbb{R}$ we have $x-1<[x]\leqslant x$, it follows that
@@
Then $v^{(n)}_{i}\geqslant\displaystyle\frac{\tau_{in}}{n}\to\tau_{i}$, $v^{(n)}_{i}\leqslant\displaystyle\frac{\tau_{i(n+3)}}{n}=\displaystyle\frac{\tau_{i(n+3)}}{n+3}\cdot\displaystyle\frac{n+3}{n}\to\tau_{i}$ when $n\to\infty$. Hence, $\nu_{i}(x)=\tau_{i}$ for all $i\in\{0,1,2,3\}$.
@@
$\hat{x}=\Delta^{4}_{\underbrace{\underbrace{0\ldots 0}_{[\tau_{01}s_{1}]}\underbrace{1\ldots 1}_{[\tau_{11}s_{1}]}\underbrace{2\ldots 2}_{[\tau_{21}s_{1}]}\underbrace{3\ldots 3}_{[\tau_{31}s_{1}]}}_{\text{1st block}}\ldots\underbrace{\underbrace{0\ldots 0}_{[\tau_{0k}s_{k}]}\underbrace{1\ldots 1}_{[\tau_{1k}s_{k}]}\underbrace{2\ldots 2}_{[\tau_{2k}s_{k}]}\underbrace{3\ldots 3}_{[\tau_{3k}s_{k}]}}_{\text{kth block}}\ldots}.$ (1)
@@
Since $\tau_{0k}s_{k}]+[\tau_{1k}s_{k}]+[\tau_{2k}s_{k}]+[\tau_{3k}s_{k}]>\tau_{0k}s_{k-1}+\tau_{1k}s_{k-1}+\tau_{2k}s_{k-1}+\tau_{3k}s_{k-1}=s_{k-4}\to\infty$ when $k\to\infty,$ then the number $\hat{x}$ is constructed correctly.
@@
then if $k\to\infty$ we have
@@
Therefore, by the Stolz theorem [15, с. 67]
@@
Let $n$ be a sufficiently large natural number, and suppose the nn–th digit of the number $\hat{x}$ falls within the kk==th block. Then

### 05 Mar 2026 v2

@@ 1. Introduction
where $(i)$ denotes the period in the ss–adic representation of the number. We call such numbers ss–adic rational. To define the nn–th digit $\alpha_{n}(x)$ of a number $x$ as a function of $x$, we agree to use only the first ss–adic representation, namely, the one that has period $(0)$.
@@
The concept of the asymptotic mean of digits of a number is closely related to the concept of digit frequency. In the case of the binary numeral system these concepts coincide. Let us recall this notion.
Let $N_{i}(x,n)$ denote the quantity of digits $i\in\mathcal{A}s$ in the ss–adic representation $\Delta^{s}{\alpha_{1}\alpha_{2}\ldots\alpha_{k}\ldots}$ of a number $x\in[0,1]$ up to and including the nn–th place, that is,
@@
We call a number $x$ normal on the base $s$ (weakly normal) if for each $i\in\mathcal{A}_{s}$ the frequency exists and equals $\nu_{i}(x)=s^{-1}$.
@@
We focus on the case $s=4$, since the case $s=3$ is analyzed in detail in our paper [10]. In fact, for $s>3$ the set $S_{\theta}$ exhibits richer properties.
@@ 2. The set $\Theta_{1}$ and the Besicovitch–Eggleston sets
If $\theta=0$ or $\theta=3$, then $\Theta_{1}$ is an anomalously fractal and everywhere dense set.
@@
Let $\theta=0$. If $\lim\limits_{n\to\infty}r_{n}(x)=0$ then for each $i\in{1,2,3}$ the following condition holds: $0\leqslant v^{(i)}_{n}(x)\leqslant v^{(1)}_{n}(x)+2v^{(2)}_{n}(x)+3v^{(3)}_{n}(x)=r_{n}(x)\to 0$ as $n\to\infty$, hence $\nu_{i}(x)=\lim\limits_{n\to\infty}v^{(i)}_{n}(x)=0$ and respectively $\nu_{0}(x)=1$. Therefore, $S_{\theta}=\Theta_{1}=E[1,0,0,0]$. This set is everywhere dense set and its Hausdorff–Besicovitch dimension is equal to
@@
Now let $\theta=3$. If $\lim\limits_{n\to\infty}r_{n}(x)=3$, then multiplying the first equation of system ()(*) by 3 and subtracting the second equation of the system, we obtain: $3v^{(0)}_{n}+2v^{(1)}_{n}+v^{(2)}_{n}=3-r_{n}$. Hence, $0\leqslant v^{(i)}_{n}(x)\leqslant 3v^{(0)}_{n}(x)+2v^{(1)}_{n}(x)+v^{(2)}_{n}(x)=3-r_{n}(x)\to 0$ as $n\to\infty$. Hence, $\nu_{i}(x)=0$ for all $i\in\{0,1,2\}$ and $\nu_{3}(x)=1$. Therefore, $\Theta_{2}=\Theta_{3}=\varnothing$ and $S_{\theta}=\Theta_{1}=E[0,0,0,1]$. This set is everywhere dense, and its Hausdorff–Besicovitch dimension equals 0. ∎
If the 44–adic representation of a number $x$ has frequencies of all digits $\nu_{0}$, $\nu_{1}$, $\nu_{2}$, $\nu_{3}$, then it has an asymptotic mean of digits $r(x)$, given by equality
@@
If $\theta\neq\displaystyle\frac{3}{2}$, then $\Theta_{1}$ contains no normal numbers. Since almost all numbers (in the sense of Lebesgue measure) are normal, then we have $\lambda(\Theta_{1})=0$. On the other hand, for $\theta=\displaystyle\frac{3}{2}$, the following inclusion holds: $\Theta_{1}\supset E\left[\displaystyle\frac{1}{4},\displaystyle\frac{1}{4},\displaystyle\frac{1}{4},\displaystyle\frac{1}{4}\right]$, where $\lambda(E)=1$. Thus, $\lambda(\Theta_{1})=1$.
@@
Since $E[p_{0},p_{1},p_{2},p_{3}]\subset\Theta_{1}$, then $\alpha_{0}(\Theta_{1})\geqslant\alpha_{0}(E[p_{0},p_{1},p_{2},p_{3}])=-\displaystyle\frac{m(\theta)}{\ln 4}.$
@@ 3. An example of a number from the set $\Theta_{1}$
We construct the number $x$ as follows. In the first step, we sequentially write $\tau^{\prime}_{01}$ zeros, $\tau^{\prime}_{11}$ ones, $\tau^{\prime}_{21}$ twos, and $\tau^{\prime}_{31}$ threes. After the $k$-th step, in the $(k+1)$-step we append to the alredy written sequence of $\tau^{\prime}_{0k}$ zeros, $\tau^{\prime}_{1k}$ ones, $\tau^{\prime}_{2k}$ twos, and $\tau^{\prime}_{3k}$ threes. As a result, among the first $\sum\limits_{i=0}^{3}\tau_{in}$ symbols of the 44–adic representation of $x$ there are exactly $\tau_{in}$ digits equal to ii. Let $n$ be a sufficiently large natural number. Since for any $x\in\mathbb{R}$ we have $x-1<[x]\leqslant x$, then it follows that
@@
Then $v^{(n)}_{i}\geqslant\displaystyle\frac{\tau_{in}}{n}\to\tau_{i}$, $v^{(n)}_{i}\leqslant\displaystyle\frac{\tau_{i(n+3)}}{n}=\displaystyle\frac{\tau_{i(n+3)}}{n+3}\cdot\displaystyle\frac{n+3}{n}\to\tau_{i}$ as $n\to\infty$. Hence, $\nu_{i}(x)=\tau_{i}$ for all $i\in\{0,1,2,3\}$.
@@
$\hat{x}=\Delta^{4}_{\underbrace{\underbrace{0\ldots 0}_{[\tau_{01}s_{1}]}\underbrace{1\ldots 1}_{[\tau_{11}s_{1}]}\underbrace{2\ldots 2}_{[\tau_{21}s_{1}]}\underbrace{3\ldots 3}_{[\tau_{31}s_{1}]}}_{\text{1st block}}\ldots\underbrace{\underbrace{0\ldots 0}_{[\tau_{0k}s_{k}]}\underbrace{1\ldots 1}_{[\tau_{1k}s_{k}]}\underbrace{2\ldots 2}_{[\tau_{2k}s_{k}]}\underbrace{3\ldots 3}_{[\tau_{3k}s_{k}]}}_{\text{k-th block}}\ldots}.$ (1)
@@
Since $\tau_{0k}s_{k}]+[\tau_{1k}s_{k}]+[\tau_{2k}s_{k}]+[\tau_{3k}s_{k}]>\tau_{0k}s_{k-1}+\tau_{1k}s_{k-1}+\tau_{2k}s_{k-1}+\tau_{3k}s_{k-1}=s_{k-4}\to\infty$ as $k\to\infty,$ then the number $\hat{x}$ is constructed correctly.
@@
then as $k\to\infty$ we have
@@
Therefore, by the Stolz theorem [15, p. 67] we have
@@
Let $n$ be a sufficiently large natural number, and suppose the nn–th digit of the number $\hat{x}$ falls within the kk–th block. Then